### 03 Dec 2023 v2

# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

Abstract
The surge in real-time data collection across various industries has underscored the need for advanced anomaly detection in both univariate and multivariate time series data. Traditional methods, while comprehensive, often struggle to capture the complex interdependencies in such data. This paper introduces TransNAS-TSAD, a novel framework that synergizes transformer architecture with neural architecture search (NAS), enhanced through NSGA-II algorithm optimization. This innovative approach effectively tackles the complexities of both univariate and multivariate time series, balancing computational efficiency with detection accuracy. Our evaluation reveals that TransNAS-TSAD surpasses conventional anomaly detection models, demonstrating marked improvements in diverse data scenarios. We also propose the Efficiency-Accuracy-Complexity Score (EACS) as a new metric for assessing model performance, emphasizing the crucial balance between accuracy and computational resources. TransNAS-TSAD sets a new benchmark in time series anomaly detection, offering a versatile, efficient solution for complex real-world applications. This research paves the way for future developments in the field, highlighting its potential in a wide range of industry applications.
keywords:
Time Series Anomaly Detection, Transformers in Machine Learning, Neural Architecture Search (NAS), NSGA-II Optimization, Deep Learning for Time Series Analysis, Adversarial Learning Paradigms
[…]
[1]organization=Department of Computer Science, addressline=University of Vermont,postcode=05405,city=Burlington, VT, country=U.S.A
1 Introduction
The prevalence of time series data in various sectors—from finance to healthcare, from infrastructure to manufacturing—highlights its critical role in modern analytics. Such data are often pivotal in uncovering underlying patterns and dependencies that are vital for strategic insights and anomaly detection, which can indicate significant shifts in system behaviors or emerging issues (Brophy et al., (2023); Li et al., (2023)).
Traditional statistical methods have long been the mainstay for anomaly detection in time series data, trusted for their robust mathematical frameworks (Thudumu et al., (2020); Wang et al., (2011); Leadbetter, (1991)). Yet, the advent of Big Data—with its hallmark features of volume, velocity, and variety—has stretched these methods to their limits, revealing shortcomings through high false positive rates and missed detections (Ashabi et al., (2020); Al-Sai et al., (2019)).
In contrast, deep learning has ushered in a paradigm shift, offering enhanced adaptability and performance in anomaly detection (Ismail Fawaz et al., (2019); Ma et al., (2021); Haq et al., (2023)). The transformer architecture, in particular, has emerged as a groundbreaking development due to its self-attention mechanism, enabling significant strides not only in natural language processing but also in time series analysis (Vaswani et al., (2017); Tuli et al., (2022); Kim et al., (2023)).
However, the specific application of transformer models to time series anomaly detection necessitates further exploration to fine-tune their structures and parameters for optimal performance (Arslan et al., (2023)). Our initiative, TransNAS-TSAD, represents a significant shift from traditional transformer models, such as TranAD, to a more dynamic and adaptable framework. This new model integrates the transformer’s architectural strengths with systematic optimization through neural architecture search (NAS), emphasizing the multi-objective optimization strengths of the NSGA-II algorithm (Elsken et al., (2019); Deb et al., (2002)).
TransNAS-TSAD marks a significant evolution from TranAD’s static architecture, primarily through the integration of NSGA-II. This shift, grounded in (Lu et al., (2023)) study, leverages NSGA-II’s proven superiority in exploring complex search spaces more efficiently than recent algorithms. By adeptly balancing performance with computational efficiency, NSGA-II enhances our neural architecture design for anomaly detection, offering improved adaptability and efficiency. (Lu et al., (2020).
1.1 Research Context and Challenges
Adapting anomaly detection models for time series data is a complex task, especially when dealing with multivariate series that present high-dimensional challenges and necessitate sophisticated model architectures (Wei, (2018); Liu et al., (2018)). The selective attention capabilities of the transformer model present a compelling solution, but only if the architectures are tailored to the specific characteristics of the dataset (Lian et al., (2019)).
The marriage of NAS with multi-objective optimization algorithms like NSGA-II represents a strategic evolution in automated model design, offering a methodical path to discovering optimal architectures that consider multiple, often conflicting, objectives (Liu et al., (2021); Xue et al., (2023)).
1.2 Theoretical Foundations and Methodological Approach
Our methodology builds on the interplay between the theoretical principles of evolutionary algorithms and the empirical strengths of deep learning architectures. We aim to craft a robust NAS framework that can intelligently navigate and select transformer architectures with hyperparameters fine-tuned for anomaly detection in time series data (Wang et al., (2021); Ma et al., (2021)). The use of NSGA-II is central to this approach, providing a mechanism to balance the trade-offs between detection accuracy and model complexity (Wang et al., (2021)).
1.3 Guiding Philosophies and Contributions
Our research is guided by the following key principles:
  • 1.
    Adaptation to Time Series Data: We assert the need for specialized models that respect the unique temporal and dimensional characteristics of time series data.
  • 2.
    Maximizing Transformer Utility: We aim to exploit the full capabilities of the transformer architecture, particularly for temporal data interpretation.
  • 3.
    Strategic Optimization via NAS: We employ NAS, especially NSGA-II, as a systematic tool for refining architectures towards optimal model configurations.
Our contributions are as follows:
  1. 1.
    Novel Application of NAS: We leverage NAS to uniquely optimize transformer architectures for the domain of time series anomaly detection.
  2. 2.
    Architectural Parameter Exploration: We conduct extensive experimentation to identify transformer configurations that redefine design norms and improve detection performance.
  3. 3.
    Benchmarking Performance: We set new standards for performance, outperforming existing methods in a variety of datasets.
  4. 4.
    Model Design Efficiency: Our approach yields models that balance computational efficiency with analytical accuracy.
  5. 5.
    Architectural and Search Strategy Synergy: We seamlessly blend the structural advantages of transformers with the adaptive capabilities of NAS, resulting in models adept for time series anomaly detection.
2 Related Work
2.1 Anomaly Detection in Time Series Data
Anomaly detection in time series data has been a focal point of research due to its critical applications across various domains such as finance, healthcare, and industrial automation. Traditional approaches have primarily relied on statistical methods, which assume specific data distributions or apply distance-based metrics (Leadbetter, (1991); Wang et al., (2011)). However, these methods often struggle with high-dimensional, complex time series data inherent in modern applications (Hodge & Austin, (2004); Patcha & Park, (2007)).
With the advent of Big Data, researchers have sought more robust solutions capable of handling large volumes of high-velocity, high-variety data. Machine learning, and specifically deep learning, has emerged as a powerful approach to detect anomalies by learning representations of normal behavior and identifying deviations (Lee et al., (2022); Landauer et al., (2023); Al-amri et al., (2021)).
2.2 Transformers in Time Series Analysis
, The transformer architecture, introduced by Vaswani et al., (2017) has become a cornerstone for sequence modeling tasks, initially within the realm of natural language, processing and subsequently in time series analysis (Li et al., (2019); Tang & Matteson, (2021). Its self-attention mechanism allows it to capture temporal dependencies without the constraints of sequential data processing, distinguishing it from its predecessors like RNNs and LSTMs (Karita et al., (2019); Reza et al., (2022); Katrompas et al., (2022)).
In the domain of time series anomaly detection, transformers have shown promise due to their ability to model complex patterns over time (Tuli et al., (2022); Liu et al., (2021); Wang et al., (2022)). However, the research is still nascent, with considerable scope for optimization and adaptation to various time series characteristics (Luo et al., (2021); Wen et al., (2022)).
2.3 Neural Architecture Search (NAS)
NAS has emerged as a paradigm to automate the design of neural network architectures. Early NAS approaches were computationally intensive, employing reinforcement learning or evolutionary algorithms to explore the design space (Liu et al., (2021); Chen et al., (2019); Chu et al., (2020)). The development of efficient NAS methods has been a recent trend, focused on reducing the computational resources required for the search process (Jin et al., (2019); He et al., (2020)).
2.4 Multi-Objective Optimization in NAS
Multi-objective optimization in NAS aims to find a trade-off between competing objectives, such as model performance and computational cost. NSGA-II, an evolutionary algorithm, has been widely used for this purpose due to its efficiency in finding a diverse set of Pareto optimal solutions (Wang et al., (2020); Lu et al., (2019); Ying et al., (2020)). The application of multi-objective optimization in NAS for time series anomaly detection is still an under-explored area, offering a fertile ground for research (Borchert et al., (2022); Lu et al., (2021)).
2.5 Convergence of NAS and Transformers
The intersection of NAS and transformer models represents the cutting edge of research in the quest for optimal architectures for specific tasks. In time series anomaly detection, this convergence promises architectures that are both performant and efficient, addressing the unique challenges of temporal data (Wen et al., (2022)).
This work builds upon these foundations, aiming to address the identified gaps by leveraging the strengths of NAS and the transformer model to push the boundaries of anomaly detection in time series data.
3 Methodology
In this section, we delineate the comprehensive methodology employed in our study, designed to effectively detect and analyze anomalies in multivariate time series data using the TransNAS-TSAD framework. We commence with a clear definition of the problem, setting the stage for the subsequent methodological steps. This is followed by an in-depth discussion on data refinement techniques, ensuring that the data is appropriately pre-processed for our analysis. Next, we delve into the intricacies of our transformer architecture, highlighting its design and capabilities. A critical component of our approach is the neural architecture search (NAS) process, employing the NSGA-II algorithm for optimizing the transformer model, which we expound upon in detail. This is supplemented by an overview of the evolutionary process that iteratively refines our model architecture. We then transition to discussing advanced anomaly detection techniques, integrating adversarial elements into our model. The section culminates with a description of how we harness the full potential of TransNAS-TSAD for practical anomaly detection applications, illustrating the real-world implications of our research. Through this structured approach, we aim to provide a clear and logical flow of our methodology, ensuring that each step contributes coherently to achieving our research objectives.
3.1 Problem Definition
Our research pioneers the application of a multi-objective approach to anomaly detection in multivariate time series data, utilizing transformer models. The problem is defined as:
Given a sequence of multivariate time series data $S=\{x_{1},x_{2},\ldots,x_{T}\}$, where each point $x_{t}$ is timestamped $t$ and resides in an $m$-dimensional space ($x_{t}\in\mathbb{R}^{m}$), our aims are:
  1. 1.
    Anomaly Detection in Time Series: Using a training time series $S$ and a test series $S^{\prime}$ of length $T^{\prime}$ with similar modalities, we seek to detect anomalies by predicting a binary sequence $Y=\{y_{1},y_{2},\ldots,y_{T^{\prime}}\}$, where $y_{t}$ identifies anomalies at timestamp $t$ in $S^{\prime}$ (1 indicates an anomaly).
  2. 2.
    Anomaly Component Analysis: Our goal extends to identifying anomalous components within each data point of the series, generating a detailed prediction sequence $Y=\{y_{1},y_{2},\ldots,y_{T^{\prime}}\}$, where each $y_{t}$ pinpoints the specific anomalous dimensions at timestamp $t$.
3.1.1 Multi-Objective Transformer Model Optimization
At the heart of our methodology is the use of NSGA-II based neural architecture search (NAS) for optimizing the transformer architecture, focusing on:
  • 1.
    Maximizing Anomaly Detection Accuracy: We aim to enhance the F1 score for precise anomaly detection.
  • 2.
    Minimizing Architectural Complexity: Concurrently, reducing the model’s size and computational demands for efficiency in practical deployment.
This dual-focused approach is a significant advancement in applying transformer models for anomaly detection.
3.1.2 Data Processing and Model Training
Prior to modeling, the time series data undergoes a preprocessing phase where normalization is applied to ensure uniformity of scale across all series components. The normalization process is formulated as follows:
$x_{t}\leftarrow\frac{x_{t}-\min(S)}{\max(S)-\min(S)+\epsilon}$
where $\min(S)$ and $\max(S)$ denote the component-wise minimum and maximum values observed in the training time series $S$, respectively. The term $\epsilon$ is a small constant introduced to prevent division by zero, ensuring numerical stability.
Subsequently, the time series is transformed into a set of overlapping windows, which serve as the input to our model. A context window of size $K$ is defined around each data point $x_{t}$, resulting in:
$W_{t}=\{x_{t-K+1},\dots,x_{t}\}$
where boundary cases are handled by replicating the first available data point to fill the window, maintaining a fixed size of $K$. This approach encapsulates the temporal dependencies inherent in time series data.
3.1.3 Anomaly Scoring and Thresholding
Our model is designed to assign an anomaly score to each context window $W_{t}$, as opposed to direct binary labels. This score is derived from the reconstruction error of the window, where $O_{t}$ denotes the reconstructed output corresponding to $W_{t}$. A dynamic threshold $D$ is established based on the distribution of the scores computed on the training data, facilitating the distinction between normal and anomalous windows. The precise mechanism for determining $D$ will be further elaborated upon in the subsequent sections, focusing on its adaptive nature and how it relates to the detection accuracy.
3.2 TransNAS-TSAD Transformer Architecture
In the TransNAS-TSAD framework, the transformer architecture is specially adapted for time series anomaly detection. This adaptation involves a unique configuration of the standard transformer model to align with the specific characteristics of time series data.
Data Augmentation Mechanisms Within the Model: TransNAS-TSAD integrates specific data augmentation techniques within the model architecture. These techniques, which can be fine-tuned during training, include:
  • 1.
    Gaussian Noise Augmentation (Optional): Adds Gaussian noise to the input sequence for variability and robustness.
  • 2.
    Time Warping & Time Masking Augmentation (Optional): Applies time warping or masking, enhancing the model’s generalization capabilities.
Encoder Operations: The encoder in TransNAS-TSAD is designed with the following operations:
  • 1.
    Linear Embedding (Optional): Transforms the input via linear mapping to a predefined dimensional space.
  • 2.
    Layer Normalization: Normalizes the sequence for stability in subsequent operations.
  • 3.
    Positional Encoding: Uses either sinusoidal or Fourier positional encoding, depending on the architecture optimized through NAS.
  • 4.
    Multi-Head Self Attention: A key feature in TransNAS-TSAD is the multi-head attention mechanism, which is aligned with the number of features in the data, inspired by the TranAD model. It employs the standard scaled-dot product attention mechanism:
    $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\text{model}}}}\right)V$ (1)
    The multi-head attention mechanism is formulated as:
    $\text{MultiHeadAtt}(Q,K,V)=\text{Concat}(H_{1},H_{2},\ldots,H_{h})$ (2)
    where $H_{i}=\text{Attention}(Q_{i},K_{i},V_{i})$.
  • 5.
    Feedforward Neural Networks (FFNs) Post-Attention: Following the attention mechanism, the sequence is processed through a series of feedforward neural networks (FFNs). Each FFN layer, consisting of a linear transformation followed by an activation function, introduces non-linearities crucial for complex pattern recognition. The configuration of these FFN layers, particularly the number of layers and the number of neurons within each layer, is dynamically determined within a predefined range in our NAS search space. This optimization, conducted using Optuna, is designed to balance anomaly detection performance with computational efficiency, ensuring that the FFN dimensions are optimally tuned for the specific demands of time series anomaly detection.
Decoder Operations: The decoder mirrors the encoder’s complexity and includes additional transformations that integrate encoder memory, apply multi-head self-attention, and utilize feedforward neural networks, culminating in a final layer for accurate reconstruction. TransNAS-TSAD can utilize either a dual sequential decoder setup or an iterative approach, depending on the optimized configuration for maximum anomaly detection accuracy.
Figure 1: Workflow of the TransNAS-TSAD Process for Time Series Anomaly Detection using Multi-Objective Neural Architecture Search (NAS) with NSGA-II Optimization.
3.3 Neural Architecture Search (NAS) Framework for TransNAS-TSAD
The NAS framework shown in Figure 1 is pivotal in achieving our primary objective of effectively detecting anomalies in multivariate time series data. By employing the NSGA-II algorithm, our NAS process systematically navigates the vast architectural landscape of transformer models. This step is essential in identifying an architecture that not only excels in anomaly detection accuracy but also aligns with the computational constraints of practical deployment scenarios. The rigorous optimization and selection of transformer architectures through NAS directly contribute to enhancing the performance and efficiency of the TransNAS-TSAD framework, thereby fulfilling our goal of developing a robust and adaptable anomaly detection system.
3.3.1 Search Space Definition
The search space outlined in Table 1 within TransNAS-TSAD represents a diverse and extensive array of potential model architectures and hyperparameters. This assortment is meticulously designed to tailor neural network configurations specifically for time-series anomaly detection:
  • 1.
    Architectural Parameters: These parameters encompass the number and types of layers in the encoder and decoder, optimized attention mechanisms for temporal data analysis, dimensions of feedforward networks, and a variety of positional encoding and normalization methods. Collectively, they enable the model to adapt its architecture for different characteristics of time-series data.
  • 2.
    Training Hyperparameters: This category includes a broad spectrum of hyperparameters, such as learning rates, batch sizes that are adaptable to varying computational resources, and dropout rates to ensure effective regularization. Additionally, the window size parameter, ranging from 10 to 30, is crucial in defining the length of the input sequence window, which directly influences how the model processes and learns from temporal patterns. Data augmentation techniques like time warping and masking are also part of the search space, enabling the model to simulate and learn from varied anomaly scenarios effectively.
The comprehensive nature of the search space in TransNAS-TSAD allows for the exploration and optimization of a wide range of models, ensuring that the most suitable architecture and hyperparameter settings are identified for effective anomaly detection in different time-series datasets.
Parameter Type Parameter Name Search Space Description
Training
Learning Rate
$1\times 10^{-5}$ to $1\times 10^{-1}$ (log scale)
Rate at which the model learns.
Dropout Rate
0.1 to 0.5
Regularization to prevent overfitting.
Batch Size
16 to 128 (step of 16)
Number of samples per training step.
Gaussian Noise
$1\times 10^{-4}$ to $1\times 10^{-1}$ (log scale)
Noise added for robustness.
Time Warping
True, False
Augmentation technique for time series.
Time Masking
True, False
Augmentation technique to mask intervals.
Window Size
10 to 30
Length of the input sequence window for the model.
Architectural
Positional Encoding Type
Sinusoidal, Fourier
Encoding type for sequence position.
Dimension Feedforward
8 to 128 (log scale)
Size of the feedforward network.
Encoder Layers
1 to 3
Number of layers in the encoder.
Decoder Layers
1 to 3
Number of layers in the decoder.
Activation Function
ReLU, Leaky ReLU, Sigmoid, Tanh
Non-linearity after each layer.
Attention Type
Scaled Dot Product
Type of attention mechanism.
Number of Attention Heads
Equal to feature dimension
Parallel attention layers.
Use Linear Embedding
True, False
Option to use a linear embedding layer.
Layer Normalization
Layer, Batch, Instance
Type of normalization used.
Self Conditioning
True, False
Conditioning strategy for the model.
Number of FFN Layers
1 to 3
Layers in the feedforward network.
Phase Type
1phase, 2phase, Iterative
Model’s reconstruction and refinement strategy.
Table 1: Neural Architecture Search Space for TransNAS-TSAD
3.3.2 TransNAS-TSAD Evaluation Strategy
Each trial within the TransNAS-TSAD NAS process entails a rigorous evaluation of models generated with varying architectural and training hyperparameters. This evaluation adopts a multi-objective approach, primarily focusing on two critical aspects: the F1 score and the number of model parameters. The F1 score serves as a key indicator of the model’s accuracy in anomaly detection, balancing precision and recall, while the number of parameters gauges the model’s architectural complexity and computational efficiency. The objective is to identify models that not only demonstrate high proficiency in accurately detecting anomalies (as reflected in a high F1 score) but also maintain a streamlined architecture (evidenced by a lower count of parameters). This dual-criteria assessment ensures the selection of models that are both effective in performance and practical in deployment, aligning with the overarching goal of achieving optimal anomaly detection with computational resourcefulness.
3.3.3 Multi-Objective Optimization with NSGA-II
TransNAS-TSAD employs the NSGA-II algorithm to balance the objectives of performance and computational efficiency. The optimization process is defined by the following equations:
$\mathbf{F}(\mathbf{x})=[f_{1}(\mathbf{x}),f_{2}(\mathbf{x})]$ (3)
where $f_{1}(\mathbf{x})$ is the F1 score and $f_{2}(\mathbf{x})$ is a computational resource metric.
Non-dominated sorting and the calculation of crowding distance are integral parts of NSGA-II:
$\text{NonDomSort}(\mathbf{X})=\bigcup_{i=1}^{k}F_{i}$ (4)
$d(\mathbf{x})=\sum_{i=1}^{k}\left(f_{i}(\mathbf{x}^{+})-f_{i}(\mathbf{x}^{-})\right)$ (5)
3.3.4 Pareto Front Exploration
The Pareto front represents the set of non-dominated solutions, providing an optimal trade-off between conflicting objectives:
$\mathcal{P}=\{\mathbf{x}\in\mathbf{X}\mid\nexists\mathbf{y}\in\mathbf{X}:\mathbf{F}(\mathbf{y})\prec\mathbf{F}(\mathbf{x})\}$ (6)
where $\prec$ indicates that $\mathbf{y}$ dominates $\mathbf{x}$.
3.3.5 Utilization of the Pareto Front
The Pareto front derived from the NAS process provides a practical framework for model selection:
  • 1.
    Interpretation and Analysis: The front is analyzed to discern architectural trade-offs, allowing stakeholders to identify models with the desired balance between accuracy and efficiency.
  • 2.
    Informed Selection: Decision-makers can select models that align with specific performance expectations and operational constraints by examining the Pareto front.
  • 3.
    Resource Allocation: Models on the Pareto front inform resource allocation, directing investments toward architectures with the most favorable trade-offs.
Algorithm 1 provides a detailed procedure that encapsulates our approach within the Optuna framework using NSGA-II (Deb et al., (2002)) for the multi-objective optimization. This algorithm elucidates the sequential steps taken from initialization to final selection, highlighting the integration of the evolutionary search process with transformer-specific architecture and hyperparameter tuning.
Input: Training dataset, validation dataset, predefined search space for transformer architectures and hyperparameters
Output: Optimized transformer architecture and hyperparameters for time series anomaly detection
1 Initialization: Set up the Optuna framework with the NSGA-II sampler. Initialize the search with a population of transformer architectures from the predefined search space. Define evolutionary process parameters: population size, number of generations, mutation and crossover rates. Specify evaluation metrics: F1 score, computational resource utilization (model size, training time), precision, recall, and inference time. for each study in Optuna do
2       for each trial in the study do
3             Define the transformer model architectural parameters and training hyperparameters within the search space constraints. Construct and train the transformer model on the Training dataset. Evaluate the model on the Validation dataset using the specified metrics. Update the study with the trial’s objectives, reflecting the model’s performance and computational efficiency. if early stopping criteria are met based on validation performance or predefined resource constraints then
4                   Terminate the trial prematurely.
5             end if
6
7       end for
8      Selection: Use Optuna’s NSGA-II sampler to select architectures for the next generation, focusing on the Pareto-optimal solutions. Crossover and Mutation: Apply Optuna’s genetic operators to evolve the study’s population, exploring new architectural variants.
9 end for
Final Selection: At the end of the NAS process, use Optuna to extract the best transformer architecture and hyperparameters from the Pareto front, considering both performance metrics and computational efficiency.
Algorithm 1 Multi-Objective NAS for Time Series Anomaly Detection with TransNAS-TSAD
3.3.6 Evolutionary Process in TransNAS-TSAD
The evolutionary process within TransNAS-TSAD serves a critical role in aligning our methodology with the overarching goal of refining anomaly detection capabilities. Through iterative refinement and selection of model architectures, this process ensures that our models are not only accurate in identifying anomalies but also evolve to become more efficient and adaptable to various data characteristics. This continuous improvement and adaptation are key to achieving high-performance anomaly detection, which stands at the core of our research objectives.
  • 1.
    Generation Loop: New models are proposed, evaluated, and bred in each generation, with superior architectures refining the Pareto front.
  • 2.
    Early Stopping: The process incorporates early stopping for models that do not show promise, thus conserving computational resources.
  • 3.
    Final Architecture Selection: The optimal architecture is selected from the Pareto front, ensuring an effective balance between the objectives.
3.4 Offline Optimization of Transformer Architecture for Anomaly Detection
Central to TransNAS-TSAD is the offline, strategic optimization of the transformer architecture, a process critical for achieving an effective balance between anomaly detection accuracy and computational efficiency. This optimization, guided by neural architecture search (NAS) using the NSGA-II algorithm, involves:
  1. 1.
    Comprehensive Trial-Based Exploration: In line with Optuna’s best practices, the process involves conducting over 100 trials to explore a wide array of transformer architectures, ensuring a thorough examination of the solution space.
  2. 2.
    Achieving Pareto Front Efficiency: The key focus of NAS is to identify and refine a collection of architectures that constitute the Pareto front, representing a spectrum of models from high-performance, resource-intensive ones to balanced and resource-efficient alternatives.
  3. 3.
    Iterative Evolutionary Process: Through evolutionary algorithms, these architectures undergo continuous adaptation and improvement, exploring innovative configurations and enhancing performance and efficiency with each generation.
3.4.1 Tailoring Architectures to Specific Deployment Needs
Post the offline NAS process, we are equipped with a diverse array of transformer architectures, each optimized for different operational contexts:
  • 1.
    For Resource-Rich Environments: High F1 score models, though computationally demanding, are suitable for scenarios where resource availability is not a constraint.
  • 2.
    In Resource-Limited Settings: The NAS process also yields architectures that are either balanced in terms of performance and resource usage or are specifically optimized for low-resource environments.
This flexibility allows for the deployment of TransNAS-TSAD in various settings, ensuring efficient and effective anomaly detection tailored to the constraints and requirements of different deployment environments.
3.5 Advanced Anomaly Detection with Adversarial Elements in TransNAS-TSAD
TransNAS-TSAD represents a significant advancement in the field of anomaly detection in time series data, strategically incorporating and extending adversarial learning paradigms. Drawing inspiration from the two-phase approach of the TranAD model, our framework elevates this concept by introducing a neural architecture search (NAS) driven methodology. This approach dynamically selects between the refined two-phase approach and our novel iterative self-adversarial approach at each trial. The integration of these advanced techniques addresses the complexities inherent in time series data, particularly in capturing subtle and sophisticated anomalies that traditional encoder-decoder architectures might overlook.
Our development of the iterative self-adversarial approach, in particular, marks a notable enhancement over existing methods. By embedding elements of adversarial training within both the two-phase and iterative approaches, TransNAS-TSAD achieves a comprehensive and adaptable anomaly detection capability. This not only enhances the sensitivity and accuracy of the detection process but also ensures robustness and adaptability across diverse datasets. The synergy of these methodologies within TransNAS-TSAD aligns seamlessly with our research goal of developing a versatile and effective solution for time series anomaly detection, demonstrating our commitment to pushing the boundaries of current technologies.
3.5.1 Two-Phase Adversarial Approach
Drawing from adversarial learning paradigms, the two-phase approach in TransNAS-TSAD seeks to enhance the model’s sensitivity to anomalies through competitive reconstruction stages.
Phase 1 - Preliminary Input Reconstruction
This phase, acting as the foundation, aims for a preliminary reconstruction of the input time-series window. The preliminary input reconstruction yields a focus score, defined by the deviation of the reconstructed output from the actual input:
$L_{\text{focus}}=\|O_{\text{initial}}-W\|^{2}$
where $O_{\text{initial}}$ is the output from the first phase and $W$ is the input time-series window. The resultant “focus score,” derived from the deviations in this initial reconstruction, serves as an attention modulator for the subsequent phase.
Phase 2 - Adversarial, Focus-Driven Reconstruction
Incorporating adversarial elements, this phase utilizes the focus scores from the phase 1 as a guiding beacon. In the adversarial phase, the second decoder aims to maximize the difference between its output and the input. Simultaneously, the first decoder aims to minimize this difference:
$L_{\text{adv1}}=\|O_{\text{adv1}}-W\|^{2}$
$L_{\text{adv2}}=-\|O_{\text{adv2}}-W\|^{2}$
where $O_{\text{adv1}}$ and $O_{\text{adv2}}$ are the outputs from the first and second decoders, respectively, during the adversarial phase.
3.5.2 Iterative Self-Adversarial Approach
Going beyond the structured two-phase reconstruction, TransNAS-TSAD introduces a dynamic iterative approach. This approach, embedded with self-adversarial mechanisms, continually refines its reconstructions.
3.5.3 Iterative Refinement:
Starting with an initial reconstruction, the deviation between the current output and the input provides feedback for the next iteration:
$L_{\text{iteration}}=\left\|O_{\text{current}}-W\right\|_{2}$
The self-adversarial mechanism can be represented as:
$L_{\text{self-adv}}=(L_{\text{iteration, prev}}-L_{\text{iteration, current}})^{2}$
where $L_{\text{iteration, prev}}$ is the loss from the previous iteration and $L_{\text{iteration, current}}$ is the loss from the current iteration. The iterative refinement continues until the change in the loss between consecutive iterations falls below a predetermined threshold, signifying convergence. Mathematically, the convergence criterion can be defined as:
$\Delta L=\left|L_{\text{iteration, current}}-L_{\text{iteration, prev}}\right|$
If $\Delta L<\epsilon$, where $\epsilon$ is a small positive value (e.g., $10^{-5}$), the iteration stops, suggesting that further refinement may not yield significant improvements. After the iterative process converges, the best anomaly score is determined. The best score is derived from the iteration with the smallest reconstruction loss, ensuring that the selected representation most closely matches the input time-series data.
3.5.4 Adaptable and Robust Anomaly Detection
What sets TransNAS-TSAD apart is its adaptability. The model is not confined to a fixed number of reconstruction phases. Depending on the intricacy of the dataset, it can dynamically adjust, ensuring that even the most subtle anomalies are not overlooked. Moreover, the incorporation of adversarial elements, both in the two-phase and iterative approaches, ensures the model continually challenges itself, leading to more refined and accurate reconstructions. TransNAS-TSAD represents a significant advancement in the realm of time-series anomaly detection. By amalgamating auto-regressive inference, adversarial mechanisms, attention modulation, and adaptive iterations, it offers a comprehensive solution, adept at detecting both overt and nuanced anomalies across varied time-series datasets.
3.6 Harnessing the Power of TransNAS-TSAD for Anomaly Detection
Building upon the foundational architecture and adversarial mechanisms of TransNAS-TSAD, we now delve into the practical realm of anomaly detection. This section elucidates the methodologies employed to infer anomalies from time-series data, leveraging the trained transformer model and various augmentation strategies.
3.6.1 The Anomaly Scoring Mechanism
Every incoming data point, denoted as $\hat{W}$, is assessed for anomalies by computing a score that quantitatively reflects its deviation from expected patterns. In this context, $\hat{W}$ represents the new window of the time series data under consideration, and $\hat{C}$ embodies any associated contextual information that could influence the anomaly detection process. The anomaly score, $s$, is calculated as:
$s=\frac{1}{2}\|R_{1}-\hat{W}\|_{2}^{2}+\frac{1}{2}\|\hat{R}_{2}-\hat{W}\|_{2}^{2}$ (7)
where $R_{1}$ and $\hat{R}_{2}$ are the reconstructed outputs from the transformer model, corresponding to different stages of the inference process.
3.6.2 Dual Inference Pathways
TransNAS-TSAD boasts of two distinct inference methodologies:
  • 1.
    Two-Phase Approach: A structured two-step inference process yielding reconstructions $(R_{1},\hat{R_{2}})$.
  • 2.
    Iterative Refinement: For scenarios demanding intricate attention, we engage in iterative inference, perfecting the anomaly scores with each iteration.
Figure 2 illustrates the anomaly detection process on dimension 0 of the SMD test dataset using TransNAS-TSAD. The upper plot displays the true data points (in blue) and the predicted values (in red), while the lower plot indicates the computed anomaly scores and their corresponding labels, offering a visual interpretation of the model’s detection capabilities in identifying anomalies.
Figure 2: Anomaly detection in TransNAS-TSAD on dimension 0 of the SMD test dataset, depicting true versus predicted values and identified anomalies
3.6.3 Dynamic Anomaly Thresholding: The Sentinel
Central to our thresholding strategy is the peaks over threshold (POT) method. At its core, POT establishes a threshold where observations surpassing this mark are deemed anomalous, allowing for effective discernment of extreme data points.
Evolutionary in nature, our modified POT (mPOT) approach ensures adaptability to the ever-changing landscape of time-series data:
$mPOT(x)=POT(x)+\alpha\times\text{recent\_deviation}(x)$ (8)
where $\alpha$ is a weight factor. The function recent_deviation calculates the deviation of the latest data points from their median value. Anomalies are pinpointed whenever any dimension’s score, $s_{i}$, exceeds this dynamic mPOT threshold.
3.6.4 Augmentative Strategies for Enhanced Precision
TransNAS-TSAD employs various augmentative strategies to bolster the precision of the POT method, ensuring that the anomaly detection mechanism remains sensitive, adaptable, and robust. These strategies aim to enhance the detection capabilities by refining the anomaly scores and the thresholds against which they are evaluated.
Moving Average Thresholding (MAT)
MAT is a dynamic thresholding technique that complements the POT method. Instead of using a static threshold, MAT calculates a moving average of recent anomaly scores to adaptively set the threshold. This dynamic adjustment ensures that the threshold is responsive to emerging data trends and patterns, enhancing its relevance and accuracy. The moving average threshold at time $t$ is given by:
$MAT(t)=\frac{1}{N}\sum_{i=t-N}^{t}s_{i}$ (9)
where $(s_{i}$ is the anomaly score at time $i$ and $N$ represents the window size for the moving average.
Rolling Statistics for Nuanced Detection
Rolling statistics, particularly the rolling mean and standard deviation, provide additional context to the POT method by capturing temporal dependencies and trends in the data. These statistics are instrumental in uncovering subtle anomalies that might otherwise be overlooked. For each data point at time $t$, the rolling mean $\mu(t)$ and rolling standard deviation $\sigma(t)$ are computed as:
$\mu(t)=\frac{1}{W}\sum_{i=t-W}^{t}x_{i}\quad\text{and}\quad\sigma(t)=\sqrt{\frac{1}{W}\sum_{i=t-W}^{t}(x_{i}-\mu(t))^{2}}$ (10)
where $x_{i}$ represents the data point at time $i$ and $W$ is the rolling window size. These rolling statistics are incorporated into the feature set, enhancing the model’s ability to discern intricate data variations and improving the relevance of the anomaly scores generated for the POT method.
4 Experimental Setup for TransNAS-TSAD
Our study rigorously evaluates the performance of TransNAS-TSAD against a suite of established benchmark models, including TranAD (Tuli et al., (2022)), LSTM-NDT (Hundman et al., (2018)), DAGMM (Zong et al., (2018)), OmniAnomaly (Su et al., (2019)), MSCRED (Zhang et al., (2019)), MAD-GAN (Li et al., (2019)), USAD (Audibert et al., (2020)), CAE-M (Zhang et al., (2021)), and GDN (Deng & Hooi, (2021)). To ensure a fair comparison, we utilize the hyperparameter configurations as specified in the original publications of these models, relying on their publicly available implementations.
The experimental infrastructure comprises a Google Colab Pro environment, leveraging a NVIDIA Tesla T4 GPU (16GB memory) and an Intel Xeon E5-2670 v3 CPU (8 cores, 51GB memory). The implementation is grounded in Python, utilizing robust libraries such as PyTorch for deep learning, Optuna for hyperparameter optimization, and pandas for data handling. These tools are instrumental in realizing the model implementation, executing the NSGA-II based NAS algorithm, and managing dataset operations.
A pivotal element of our methodology is the neural architecture search (NAS), which automates the architectural design, navigating a comprehensive search space of architectural and training hyperparameters (detailed in Table 1). This search space is meticulously crafted to enable the NAS algorithm to adapt the model architecture to the unique characteristics of each dataset.
Consistent with the practices in OmniAnomaly and TranAD, we apply the peaks over threshold (POT) method (Leadbetter, (1991)) across all datasets with a uniform coefficient of $10^{-4}$. The low quantile parameter, however, is fine-tuned for each specific dataset to maintain alignment with established benchmarks, thus ensuring an equitable basis for comparative analysis.
4.1 Datasets
Dataset Train Size Test Size Dimensions Sequences Anomalies (%)
NAB 4,033 4,033 1 6 0.92
UCR 1,600 5,900 1 4 1.88
MBA 100,000 100,000 2 8 0.14
SMAP 135,183 427,617 25 55 13.13
MSL 58,317 73,729 55 3 10.72
SWaT 496,800 449,919 51 1 11.98
WADI 1,048,571 172,801 123 1 5.99
SMD 708,405 708,420 38 4 4.16
Table 2: Dataset Statistics
Our study employs a comprehensive suite of datasets, each meticulously chosen to challenge and validate the robustness of time series anomaly detection methods across a variety of real-world scenarios. Table 2 summarizes the statistics of the datasets. These datasets, renowned for their complexity and diversity, allow for an effective benchmarking of our TransNAS-TSAD approach against existing methodologies.
  • 1.
    Numenta Anomaly Benchmark (NAB) Dataset: OThis dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City. It’s important to note some anomalies in labeling within this dataset, particularly in the nyc-taxi trace, which is often excluded from analyses (Ahmad et al., (2017); Nakamura et al., (2020)).
  • 2.
    HexagonML (UCR) Dataset: Featured in the KDD 2021 cup, this dataset comprises a diverse collection of univariate time series. In our research, we selectively utilized datasets that are derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets. This choice was driven by our need for natural data representations in our analysis. Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) Dataset (Dau et al., (2018)).”
  • 3.
    MIT-BIH Supraventricular Arrhythmia Database (MBA) Dataset: This dataset includes electrocardiogram recordings from four patients, featuring anomalies such as supraventricular contractions or premature heartbeats. It’s a well-recognized dataset in data management studies (Moody & Mark, (2001); Boniol et al., (2020)).
  • 4.
    Soil Moisture Active Passive (SMAP) Dataset:: Provided by NASA, this dataset consists of soil samples and telemetry data from the Mars rover (Hundman et al., (2018)).
  • 5.
    Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.(Hundman et al., (2018)) Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by (Nakamura et al., (2020); Tuli et al., (2022)).
  • 6.
    Secure Water Treatment (SWaT) Dataset: This dataset is derived from a real-world water treatment plant, including data from 7 days of normal operations and 4 days under abnormal conditions. It features readings from various sensors and actuators (Mathur & Tippenhauer, (2016)).
  • 7.
    Water Distribution (WADI) Dataset: An expansion of the SWaT dataset, WADI includes a larger array of sensors and actuators. The dataset spans over a longer period, with 14 days of normal operation and 2 days under attack scenarios (Ahmed et al., (2017)).
  • 8.
    Server Machine Dataset (SMD): Covering five weeks of data, this dataset includes resource utilization traces from 28 machines in a compute cluster. Only specific non-trivial sequences are used for analysis (Su et al., (2019)).
Each dataset is selected for its potential to challenge and validate the robustness of time series anomaly detection methods across various industry scenarios.
4.2 Evaluation Criteria
In our study, aimed at advancing anomaly detection in time series data, we employ key metrics aligned with the objectives of the TransNAS-TSAD framework. These metrics are selected to optimize anomaly detection effectiveness, considering practical deployment aspects.
4.2.1 F1 Score as Primary Performance Indicator
The F1 score, the harmonic mean of precision and recall, is central to our evaluation strategy. It is particularly effective in addressing the challenge of imbalanced datasets common in anomaly detection, offering a balanced performance measure by accounting for the trade-off between false positives and false negatives.
4.2.2 Introducing the Efficiency-Accuracy-Complexity Score (EACS)
Model complexity is often indicated by parameter count. However, this metric can be misleading, especially for transformer-based models like TransNAS-TSAD. These models, despite their high parameter count, capture complex data patterns effectively, justifying their complexity. Therefore, we report parameter count and interpret it in the context of the model’s advanced capabilities. To comprehensively evaluate TransNAS-TSAD, we introduce the Efficiency-Accuracy-Complexity Score (EACS), a composite metric encompassing accuracy, training efficiency, and model complexity. It is defined as:
$EACS=w_{\text{acc}}\times\left(\frac{\text{F1 Score}}{\text{Max F1 Score}}\right)+w_{\text{time}}\times\left(\frac{\text{Training Time}}{\text{Max Training Time}}\right)+w_{\text{params}}\times\left(\frac{\text{Parameter Count}}{\text{Max Parameter Count}}\right)$
where $w_{acc}$, $w_{time}$, and $w_{params}$ are weights for the normalized F1 score, inverse training time, and inverse parameter count, respectively. Here, the weights are set as $w_{acc}=0.4$, $w_{time}=0.4$, and $w_{params}=0.2$, balancing accuracy, training efficiency, and model simplicity. This weighting scheme aligns with our research goals and practical deployment considerations, prioritizing accuracy and training efficiency while considering model complexity.
Transformer models, known for their high parameter count, offer efficient processing capabilities through parallelization, making them suitable for scenarios requiring high accuracy where computational resources are not a constraint.
4.3 Note on Additional Metrics
Other metrics, such as ROC/AUC, are also valuable; however, our focus on the F1 score, parameter count, and EACS aligns with the specific goals of TransNAS-TSAD. This approach ensures a balanced evaluation of the models, considering both their effectiveness in anomaly detection and practical deployability.
5 Results
5.1 Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
In our detailed evaluation of anomaly detection models, as captured in Table 3, the TransNAS-TSAD model emerges as a formidable contender, showcasing its prowess across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, with metrics of precision, recall, and F1 scores serving as the benchmarks.
Standing out for its exceptional performance, TransNAS-TSAD notably surpasses other models in datasets such as NAB, UCR, MBA, and SMAP, with remarkable F1 scores (0.9411, 0.9910, 0.9861, and 0.9510 respectively), thereby demonstrating its adeptness in handling both univariate and multivariate time series. This is particularly evident when compared to models like OmniAnomaly and MSCRED, where TransNAS-TSAD’s ability to navigate complex data patterns becomes apparent.
However, in the MSL dataset, the TranAD model, with its nuanced approach, shows a marginally superior performance with an F1 score of 0.9494 compared to TransNAS-TSAD’s 0.9482, possibly indicative of TranAD’s specific strengths in handling the dataset’s characteristics. Despite this, TransNAS-TSAD regains its lead in the SWaT dataset with an F1 score of 0.8314, highlighting its adaptability to various industrial contexts.
Its exemplary performance continues in the challenging WADI and SMD datasets, where it achieves the highest F1 scores of 0.7735 and 0.9986, respectively, showcasing its unrivaled efficiency in detecting subtle anomalies in complex, high-dimensional datasets. This overarching success of TransNAS-TSAD, especially when contrasted with TranAD’s performance, underscores its comprehensive capability in anomaly detection.
While models like LSTM-NDT and DAGMM show variable performance across different datasets, their inconsistency underlines TransNAS-TSAD’s steady effectiveness across various scenarios. Similarly, while OmniAnomaly and MSCRED perform well in certain contexts, they fall short of TransNAS-TSAD’s sophisticated temporal pattern handling.
Other competitive models, including MAD-GAN and USAD, are eclipsed by TransNAS-TSAD’s more holistic approach, and even though CAE-M and GDN demonstrate strengths in specific scenarios, they generally do not match TransNAS-TSAD’s overall adaptability.
The superiority of TransNAS-TSAD, especially in comparison to TranAD, invites further investigation through critical difference analysis, which would reinforce the statistical significance of the observed performance differences and provide a deeper understanding of TransNAS-TSAD’s robustness and reliability against a backdrop of diverse and challenging anomaly detection scenarios.
Method NAB UCR MBA SMAP
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879
DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888
OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728
MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664
MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654
USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419
CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827
GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518
TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915
TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510
Method MSL SWaT WADI SMD
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042
DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491
OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401
MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414
MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915
USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495
CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367
GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342
TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605
TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8626 0.7340 0.7735 0.9985 0.9988 0.9986
Table 3: Benchmarking TransNAS-TSAD against Baseline Models: A summary of precision (P), recall (R), and F1 scores showcases the advanced anomaly detection capabilities of TransNAS-TSAD, derived via NSGA-II-based NAS, across diverse datasets.
5.2 In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
Furthering our investigation from Table 3, Table 4 provides a more granular look at the training efficiency and overall performance of anomaly detection models, with a particular emphasis on the Efficiency-Accuracy-Complexity Score (EACS). This table not only reiterates the superior performance of TransNAS-TSAD in precision and recall metrics, as observed in Table 3 but also introduces a critical analysis of training times and parameter counts across a range of datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD.
The training time, a pivotal yet often overlooked aspect, emerges as a decisive factor in evaluating the practicality of models. In operational settings, where rapid model deployment and iterative updating are essential, training time gains precedence over traditional metrics like parameter count. This is evidenced in the case of TransNAS-TSAD, which, across diverse datasets, demonstrates not only high F1 scores but also markedly reduced training times, leading to its high EACS. In the NAB dataset, for instance, TransNAS-TSAD achieves an F1 score of 94.11% with a training time of just 2.70 sec, resulting in an EACS of 0.9742. This trend of balancing accuracy with training efficiency is replicated in the UCR dataset, where it achieves an F1 score of 99.10% and an even lower training time of 2.24 sec, leading to an EACS of 0.9922.
Diving deeper, in datasets like MBA and SMAP, TransNAS-TSAD’s EACS scores of 0.9932 and 0.9734, respectively, are not solely the product of its accuracy but are significantly bolstered by its training time efficiency. For example, in MBA, its training time of 2.27 is among the lowest, demonstrating the model’s capability to quickly adapt and learn from complex multivariate time series data. In comparison, models such as MSCRED, while exhibiting commendable F1 scores, are penalized in their EACS due to their disproportionately high training times (e.g., 774.99 sec for MBA), highlighting a crucial trade-off between model complexity and operational efficiency.
In more challenging datasets like SWaT and WADI, the importance of training time becomes even more pronounced. TransNAS-TSAD’s ability to maintain high F1 scores while drastically cutting down on training times positions it as a highly effective tool for real-world applications. Its EACS of 0.9297 for SWaT and 0.6977 for WADI, while highlighting its proficiency in these industrial datasets, also reflect the balanced approach of the model, juxtaposing accuracy with swift adaptability.
Furthermore, the analysis of models like GDN and TranAD provides additional context. While GDN shows strong performance in specific datasets like MSL with the highest F1 score (95.91%) and a respectable EACS of 0.9191, it falls behind TransNAS-TSAD in terms of training efficiency. Similarly, TranAD, despite its high F1 scores and relatively lower parameter counts, does not consistently match the training efficiency of TransNAS-TSAD, as seen in datasets like SMD and WADI.
In conclusion, Table 4’s comprehensive analysis underscores the critical role of training time in the practical deployment of anomaly detection models. It establishes that models with lower training times, like TransNAS-TSAD, offer a significant advantage in real-world scenarios, effectively balancing accuracy, efficiency, and complexity. This detailed examination not only affirms the superiority of TransNAS-TSAD in standard performance metrics but also in its operational readiness, making it an exemplary model for efficient and effective anomaly detection in diverse and demanding environments.
Method NAB UCR MBA SMAP
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459
DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050
OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230
MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982
MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683
USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439
CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529
GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440
TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445
TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734
Method MSL SWaT WADI SMD
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5737 90.42 460.70 42853 0.9100
DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6392 94.91 337.90 10516 0.9420
OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7471 94.01 311.20 22941 0.9412
MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6435 84.14 349.77 14237356 0.6978
MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7228 91.50 424.60 9903 0.9188
USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.6983 94.95 328.77 10609 0.9432
CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3203 93.67 3606.90 114219 0.5731
GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4245 83.42 1000.40 7255 0.8226
TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6316 96.05 56.70 135110 0.9760
TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 77.35 225.60 1752919 0.6977 99.81 52.40 132050 0.9916
Table 4: Comparative analysis of the Efficiency-Accuracy-Complexity Score (EACS) across various anomaly detection models, highlighting the performance of TransNAS-TSAD. The EACS is calculated by normalizing and weighting F1 scores (0.4), training times (0.4), and parameter counts (0.2), demonstrating the balanced efficiency and accuracy of TransNAS-TSAD against benchmark models across multiple datasets
Arch and Training Params SMAP UCR MBA SWaT MSL SMD NAB WADI
Learning Rate 0.0002128 0.0019997 0.003441 0.0000403 0.006192 0.0002273 0.006924 0.0000131
Dropout Rate 0.2353 0.4474 0.1795 0.3836 0.3730 0.1554 0.4560 0.2237
Dim Feedforward 101 124 41 42 121 86 10 16
Batch Size 32 48 16 32 32 48 96 112
Encoder Layers 2 1 3 2 2 3 1 2
Decoder Layers 1 2 1 3 3 1 2 1
Activation Func. sigmoid leaky_relu tanh tanh leaky_relu sigmoid relu sigmoid
Time Warping False True True True True False False False
Time Masking True True False True True False True False
Gaussian Noise 0.000151 0.027956 0.007925 0.058019 0.000428 0.000628 0.000119 0.000194
Linear Embedding False True False False True True True True
Phase Type 2phase iterative iterative 2phase 2phase 2phase iterative 2Phase
Self Conditioning False True False True True False True True
Layer Norm False False False True True True False True
Pos. Enc. Type sinusoidal fourier fourier sinusoidal fourier sinusoidal sinusoidal fourier
FFN Layers 1 1 1 3 2 3 1 3
Attn Heads 25 1 2 51 55 38 1 127
Window Size 10 20 14 22 26 12 18 26
Table 5: TransNAS-TSAD Hyperparameter Optimization Results: Detailed summary of the optimal training and architectural hyperparameters identified for each dataset, demonstrating the adaptive precision of TransNAS-TSAD in time series anomaly detection.
5.2.1 Optimal Model Configurations Identified by TransNAS-TSAD
Table 5 showcases the culmination of TransNAS-TSAD’s NSGA-II optimization process, highlighting the training and architectural hyperparameters of models that achieved the best balance between high F1 scores and low parameter counts. These configurations represent the pinnacle of the framework’s capability to identify models uniquely suited to each dataset’s anomaly detection needs.
A key observation from the table is the diversity in architectural parameters across datasets, reflecting TransNAS-TSAD’s adeptness in tailoring models to specific data characteristics. For example, in the SMAP dataset, the model employs a relatively simple architecture with fewer encoder and decoder layers, which is effective in handling the environmental time series data of this dataset. In contrast, the WADI dataset, known for its complex sensor network, necessitates a more intricate model structure, evident in its higher number of attention heads and the use of a two-phase reconstruction and refinement strategy.
Similarly, the variation in window sizes, ranging from 12 data points for SMD to 30 data points for WADI, underscores the model’s flexibility in adapting to the temporal scale of different datasets. Larger window sizes in datasets like WADI allow the model to capture longer-term dependencies and subtle anomalies in extensive time series data, a crucial requirement for sophisticated water-related infrastructures.
The adaptability of TransNAS-TSAD is further demonstrated in its choice of positional encoding and layer normalization techniques, which vary significantly among datasets. For instance, the sinusoidal positional encoding in SMAP and NAB caters to their unique temporal patterns, while the Fourier positional encoding in UCR and WADI aligns with the datasets’ complex spectral characteristics.
These results from TransNAS-TSAD’s NAS process validate the framework’s ability to not only fine-tune hyperparameters for operational efficiency but also to intricately adapt its architectural design to the nuanced demands of various time series anomaly detection scenarios, ensuring optimal detection performance across a broad spectrum of datasets.
Figure 3: Analysis of hyperparameter importance for F1 score optimization in TransNAS-TSAD across four datasets (NAB, MBA, SMAP, and WADI). The plots illustrate the relative impact of different hyperparameters on the F1 score, guiding the model’s fine-tuning process for effective anomaly detection in diverse time series datasets
5.3 Architectural and Training Hyperparameter Importance for F1 Score Optimization
The optimization of F1 scores in the TransNAS-TSAD model shown in Figure 3 is critically analyzed through hyperparameter importance plots for four datasets: NAB, MBA, SMAP, and WADI. These plots offer quantitative insights into the relative impact of various hyperparameters on the F1 score, aiding the fine-tuning process in our NAS-driven anomaly detection model.
For the NAB dataset, gaussian_noise and activation_function are the most influential hyperparameters, with importances of 0.18 and 0.15, respectively. The incorporation of gaussian_noise enhances the model’s generalization capabilities, essential for robustness, while the choice of activation_function is key for capturing the data’s nonlinear relationships.
The MBA dataset’s analysis underscores the importance of activation_function as 0.11, num_feedforward_layers as 0.07, and dropout_rate as 0.07, indicating the need for a complex model architecture to effectively learn from multivariate ECG time series.
In the SMAP dataset, gaussian_noise with importance 0.53 is paramount, highlighting the model’s need to handle noisy environmental data effectively. Additionally, batch_size with importance 0.07 influences the model’s performance, reflecting the impact of batch processing on training.
The WADI dataset, with its extensive sensor network, prioritizes activation_function with importance 0.16 and dropout_rate with importance 0.08 to maintain model robustness and prevent overfitting in high-dimensional spaces.
The recurrence of activation_function and dropout_rate across datasets emphasizes their role in non-linear data transformation and regularization. This pattern reflects the impact of data quality and structure on the model’s learning efficacy.
The conducted hyperparameter analysis illuminates the tailored performance of TransNAS-TSAD across varied datasets, with each dataset’s specifics dictating the importance of particular hyperparameters. This not only affirms the model’s adaptability but also its capacity for ongoing enhancement, ensuring it remains adept at confronting the evolving complexities of time series data.
Figure 4: Pareto front plots illustrating the trade-off between F1 score and number of parameters for NAB, MBA, SMAP, and WADI datasets in the TransNAS-TSAD optimization process.
5.4 Pareto Front Analysis for Model Optimization
Building upon the insights from the hyperparameter importance analysis, the Pareto front optimization results for the TransNAS-TSAD model, as depicted in Figure 4, further elucidate the model’s performance. The Pareto front plots for the NAB, MBA, SMAP, and WADI datasets underscore the delicate balance between model complexity, as measured by the number of parameters, and the model’s effectiveness, as quantified by the F1 score.
For the NAB dataset, the Pareto front indicates a dense congregation of model configurations. A significant number of these configurations exhibit a high parameter count without a corresponding increase in the F1 score, suggesting a potential plateau in performance gains relative to complexity. This observation prompts a critical evaluation of model parsimony, encouraging the selection of simpler models that maintain performance while mitigating the risk of overfitting.
In the MBA dataset, the dispersion of trials across the Pareto front reflects a comprehensive exploration of the architectural space. Interestingly, several models achieve commendable F1 scores without a proportionate surge in parameters, highlighting efficient architectural choices that capture the essential dynamics of ECG time series data without unnecessary complexity.
The SMAP dataset presents an outlier model with a substantial number of parameters yet achieving a high F1 score. This result may point to an overfitting scenario where the model’s complexity does not translate into generalized performance. Such an insight is invaluable for guiding the model selection process towards architectures that balance accuracy with generalization.
The Pareto front for the WADI dataset, characterized by its complex sensor network, illustrates the necessity of sophisticated models to enhance the F1 score. The trend of increasing model complexity to achieve incremental improvements in performance is evident, underlining the inherent challenges of anomaly detection in high-dimensional industrial control systems.
These Pareto fronts provide a visual and quantitative tool for identifying optimal model configurations that achieve a balance between performance and complexity. They serve as a decision-making aid for selecting models that not only excel in accuracy but are also efficient and scalable, aligning with the practical demands of anomaly detection in diverse environments.
The findings from the Pareto front analysis are integral to the ongoing development of the TransNAS-TSAD model within the broader scope of our research. They contribute to the understanding of how different model architectures perform across various datasets, offering a pathway to enhancing the model’s adaptability and ensuring its continued efficacy in the face of evolving data challenges.
6 Discussion: Addressing Challenges and Exploring Future Directions in TransNAS-TSAD
TransNAS-TSAD represents a significant advancement in anomaly detection within time series data, primarily through its innovative integration of advanced adversarial learning paradigms, NSGA-II optimization, and transformer architecture optimization. This confluence of technologies marks a substantial leap forward, particularly in the detection of subtle and complex anomalies that have typically eluded traditional methods.
A pivotal challenge addressed by TransNAS-TSAD is the balance of detection accuracy against computational efficiency. The NSGA-II algorithm plays a crucial role in this regard, effectively managing the equilibrium to ensure optimal model performance, without incurring excessive computational demands. This aspect is particularly crucial in real-world applications where resources are finite and efficiency is paramount.
The adaptability of TransNAS-TSAD is further exemplified in its dynamic adjustment capabilities, allowing it to effectively respond to the unique characteristics of different datasets. This adaptability is essential in the domain of anomaly detection, where dataset variability can present diverse challenges. Additionally, the iterative self-adversarial approach employed by TransNAS-TSAD significantly enhances detection accuracy, showcasing the model’s sophisticated capabilities in identifying anomalies.
However, challenges remain, particularly in the realm of ensuring model generalization across a diverse array of datasets. The Pareto front analysis within TransNAS-TSAD has highlighted a delicate balance between model complexity and effectiveness. Some configurations risk overfitting, which is a pertinent issue for future research endeavors. Improving generalization capabilities, without compromising detection accuracy, remains a key area for further investigation.
Looking ahead, several promising avenues for enhancement and innovation present themselves. Enhanced real-time data processing capabilities, particularly for applications in environmental monitoring and industrial control systems, represent a significant area for advancement. Techniques from data assimilation and online learning could be effectively integrated into TransNAS-TSAD to address these challenges. Additionally, the development and implementation of dynamic thresholding strategies, such as the Moving Average Thresholding (MAT) and the incorporation of rolling statistics, offer exciting prospects. These augmentative strategies could pave the way for even more nuanced detection capabilities within the TransNAS-TSAD framework.
Furthermore, the exploration of hybrid systems that synergize machine learning with simulation approaches, along with advancements in neuro-symbolic systems, could substantially enhance the model’s adaptability and effectiveness across various scenarios. Finally, a human-centric approach to machine learning, integrating human feedback in a more intuitive and formalized manner, remains a significant challenge. TransNAS-TSAD stands to benefit greatly from such integration, enhancing not only the interpretability but also the overall usability of the model in real-world applications.
As the landscape of time series anomaly detection continues to evolve, so too will the strategies and methodologies embodied in TransNAS-TSAD, ensuring its continued relevance and efficacy in this dynamic field.
7 Conclusion
TransNAS-TSAD emerges as a groundbreaking framework in the realm of time series anomaly detection, effectively integrating transformer architecture with neural architecture search and NSGA-II optimization. The model’s superior performance, evidenced across diverse datasets, demonstrates its robustness in accurately detecting anomalies in both univariate and multivariate time series data.
The strategic application of NSGA-II within TransNAS-TSAD skillfully addresses the challenge of balancing accuracy with computational efficiency, a critical consideration in practical applications. Moreover, the incorporation of advanced adversarial learning paradigms contributes to its ability to detect nuanced anomalies, a notable advancement over traditional methods.
TransNAS-TSAD’s development signifies a notable contribution to the field of anomaly detection, enhancing the capabilities and efficiency of current methodologies. As the landscape of data analysis continues to evolve, the principles and approaches embodied in TransNAS-TSAD will undoubtedly inspire further advancements in machine learning, paving the way for innovative applications in various sectors. The framework not only sets a new benchmark in anomaly detection but also opens up exciting prospects for future research, particularly in enhancing real-time processing and integrating human-centric approaches to machine learning.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. We would like to extend our gratitude to the providers of the data sets used in this work, whose contributions were invaluable to our research. The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance.